Calculus II, Spring 2019 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 02

1. Section 10.4: Problems 17, 27, 29, 31, 43, 45, 51, 61, 62.

Note in problem 62: the case p = 1, q < −1 should be convergent, instead.

2. Section 10.5: Odd numbered problems in problem 17-43, 61, 65.

3. (Optional) Remark on Example (c’) in page 12 of mse19s 190226.pdf. This problem is actually beyond this course. For those who are interested, take the following Stirling’s formula for granted

lim_{n→∞} \frac{e^{n}n!}{n^{n}\sqrt{n}} = \sqrt{2π}  (1)

and use it to find a suitable series to compare with.